At what temperature will the RMS velocity of a gas be tripled its value at STP ?

Grade:12th

1 Answers

Eshan
5 years ago
Root mean square velocity\propto \sqrt{T}

Hence,\dfrac{T'}{T_{stp}}=\dfrac{v'^2}{v_{rms}^2}=9

Therefore at 9 times the standard temperature will the rms velocity be tripled.